OpenStudy (anonymous):

ues the method of substitution to solve the system of linear equations. 2x-y=6 7x-3y=23

OpenStudy (anonymous):

multiply eqn #1 by -3--> you have -6x+3y=-18. Then add the eqn 2 and the new eqn you get x=5. from either of the original eqns, you can get y.

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

if we multiply first equation with -3 \[2x-y=6 --> -6x+3y=-18\] Now we are going to add the -6x+3y=-18 dude with 7x-3y=23. You will then have x= 23-18=5. you know that 2x-y=6, and you also know x=5. So you can find y to be 4.

OpenStudy (anonymous):

you lost me